# EE3P3D: Event-based Estimation of Periodic Phenomena Frequency using 3D Correlation

### v1

@@ Abstract
We present a novel method for measuring the frequency of periodic phenomena, e.g., rotation, flicker and vibration, by an event camera, a device asynchronously reporting brightness changes at independently operating pixels with high temporal resolution. The approach assumes that for a periodic phenomenon, a highly similar set of events is generated within a specific spatio-temporal window at a time difference corresponding to the phenomenon’s period. The sets of similar events are detected by 3D spatio-temporal correlation in the event stream space. The proposed method, EE3P3D, is evaluated on a dataset of $12$ sequences of periodic phenomena, i.e. flashing light and vibration, and periodic motion, e.g., rotation, ranging from $3.2$ Hz to 2 kHz (equivalent to 192192192192120 000120000120\,000120 000 RPM). EE3P3D significantly outperforms published methods on this dataset, achieving the mean relative error of $0.1\%$.
@@
Event camera Frequency estimation Periodic phenomena Frequency of periodic phenomena.
@@ 1 Introduction
Traditional methods for measuring periodic phenomena often involve contact-based devices such as tachometers and rotary encoders. These approaches, while established, have inherent limitations: (i) physical contact with the target object can influence its movement and introduce measurement inaccuracies, (ii) contact methods may not be feasible in scenarios with delicate objects, confined spaces, or situations where interfering with the target significantly affects its movement. Laser tachometers offer a less invasive and highly accurate [^1] See https://meters.uni-trend.com/product/ut370-series/#Specifications. alternative but require reflective material (e.g., a sticker) on the target and precise aiming of the laser, as missing the reflective material pass-through results in an inaccurate measurement. Although contactless event-based methods exist, they are either handcrafted for a specific type of periodic phenomena [2] or require user supervision [13, 14].
Figure 1: EE3P3D: (i) data captured from an event camera is aggregated into a 3D array, (ii) the array is split into same-sized areas, and in each area a template depth is automatically selected, (iii) a 3D correlation of the template with the event stream is computed, (iv) a frequency is calculated from the median of time deltas measured between correlation peaks for each window, (v) the output frequency is computed as a median of measurements from all windows.
In this paper, we propose a non-contact method for measuring the properties of any periodic phenomena using an event camera. The input of our method is the event stream, and the output is a single scalar – the frequency of the periodic phenomena. Our approach assumes that similar sets of pixel activations of an event camera (events) will occur at periodic time intervals in any sequence capturing a periodic phenomenon. The periodic intervals correspond to the target frequency.
We evaluated our method on a dataset of twelve sequences of four types of periodic phenomena: (i) light flickering, (ii) object vibration, (iii) rotational speed, and (iv) periodic movement and achieved a mean relative error of $0.1\%$, which is 140$\approx$ 140$\times$ 140 × lower than the mean relative error achieved by the best available published method (see Tab. 4).
@@ 2 Related work
Commercially available solutions for rotation speed measurement include contact and contactless methods. Contact-based devices include traditional mechanical tachometers that connect directly to a rotating shaft. These can introduce potential inaccuracies due to added mass and friction. Electrostatic sensors offer a contactless approach, detecting changes in the electromagnetic field around a rotating object (equipped with a bearing). The rotational speed is the frequency of these changes [11]. Optical encoder tachometers use a photoelectric sensor to detect light passing through a rotating disc with patterned segments. The rotation speed is the frequency of light intensity changes [1]. Finally, laser tachometers utilise a laser beam that bounces off reflective markers on the rotating target. The rotation speed is the frequency of the light reflections detected by a sensor.
@@
3 Methods
In this section, we describe three proposed methods. The Simple (Sec. 3.1) and FFT (Sec. 3.2) methods serve as baselines, and the proposed method is presented in Sec. 3.3.
3.1 A Simple baseline
A baseline method was implemented to test the hypothesis that separate per-pixel events analysis is insufficient for accurate frequency estimation of complex periodic phenomena. The event camera produces a sequence of tuples $(x,y,p,t)$, where $x,y$ are spatial coordinates, $p$ is event polarity (positive – $1$ or negative – 00), and $t$ is a timestamp in microseconds of the event creation.
For each pixel at position $x,y$, we construct a list of timestamps $T_{x,y,p}$ of a given polarity $p$.
@@ 3.1 A Simple baseline
    Frequency prediction – the final predicted frequency for each window is derived from the estimated period $T$ in microseconds using the following equation $\nu(T)={10^{6}}/{T}$, where $\nu(T)$ returns frequency in Hertz (Hz).
3.2 A FFT baseline
The FFT baseline method estimates per-pixel frequency in four steps:
@@ 3.2 A FFT baseline
    FFT computation – for every pixel and selected polarity, we compute a one-dimensional n-point Discrete Fourier Transform with the efficient Fast Fourier Transform (FFT) algorithm [3] with the timestamp of event creation represented as a Dirac pulse.
@@
    Median filtering – a 2D median filter using a local window size of $3$ is applied to the 2D array of estimated per-pixel frequencies to mitigate outliers.
@@
    Results aggregation – the mode of detected frequencies is returned.
3.3 EE3P3D - The proposed method
Broadly speaking, the event stream is quantized and 3D correlation responses analysed to detect periodic similarities along the time axis of the event data. An overview of the method is shown in Fig. 1. We identify the period of a spatio-temporal pattern in the event stream in four steps:
@@ 3.3 EE3P3D - The proposed method
    3D correlation – a 3D correlation of the template $\mathbf{T}_{i}$ is computed with the area $\mathbf{A}_{i}$ (see Fig. 2). The time difference between consecutive peaks in correlation responses is the estimated period of the phenomenon.
@@ 4 Experiments
In this section, we discuss the event camera used and the selection of parameters for methods presented in the previous section, followed by experiments measuring rotation speed, frequency of periodic light flashes, vibrations, and moving objects.
In our experiments, we used the Prophesee EVK4 HD event camera with $1280\times 720$ resolution, capturing up to $1066$ million events per second [6] under light conditions ranging from $0.08$ up to 100 000100000100\,000100 000 lux [^2] Spec sheet available at https://bit.ly/Sony-Prophesee-IMX636-IMX637-specsheet. with microsecond time resolution (equivalent to 10 0001000010\,00010 000 frames per second).
@@
We evaluated the accuracy of the Simple baseline method (see Sec. 3.1) when using window sizes $W$ of $30\times 30$ px, $45\times 45$ px, $60\times 60$ px and $75\times 75$ px and achieved the lowest mean relative error (MRE) by setting $W$ to $45\times 45$ px. Further increasing or decreasing window size or using mode instead of the median in the temporal or spatial analysis steps of this method negatively impacted the accuracy.
The FFT baseline method (see Sec. 3.2) achieved lower mean relative error when using timestamps of events with negative polarity (MRE of $60\%$ as opposed to $67.467\%$).
We evaluated all different configurations of parameters (window sizes $W$ of $30$, $45$, $60$ and $75$ and counts of template events $N$ of $100$, $200$, …, $3000$) for the proposed method (see Sec. 3.3). In all our experiments, the length of the quantisation window $t_{\text{quant}}$ is set to $100$ microseconds. When using window size $W$ set to $45\times 45$ px the method achieves the lowest MRE by a substantial margin and a low maximal relative error compared to results when using other window sizes (see Tab. 1). We noticed a significant improvement in accuracy when using counts of template events $N$ higher than $1500$ (see Tab. 2). The best accuracy was achieved using the window size $W$ set to $45\times 45$ px and the template event count $N$ set to $1800$ events. These parameters are proposed as default parameters of the method.
To objectively compare the results of the proposed method, we use the mentioned baseline methods (see Sec. 3.1, 3.2), a vibration estimation method [14] developed by Prophesee, manufacturer of the used event camera, the open-source Frequency-cam method [13] and re-implemented EB-ASM method [2]. As the output of Frequency-cam is in the form of colourmaps, we modified it to instead output a reading for each pixel that would be used for colourmap creation and computed their median. Since every method produces a different amount of measurements per one second of input data, we calculate the median of measurements to obtain a single value for a comparison with other methods.
Window size (px) $30\times 30$ $45\times 45$ $60\times 60$ $75\times 75$
Avg rel error $37.78$ $\mathbf{19.34}$ $42.96$ $487.98$
Max rel error $\mathbf{3225.00}$ $3362.50$ $7712.50$ $22737.50$
Table 1: EE3P3D - dependence on window size. Average and maximum relative errors (%) on the 12 sequence dataset.
Template event count $100-1400$ $1500-3000$
Average relative error $308.46$ $\mathbf{5.75}$
Maximum relative error $22737.50$ $\mathbf{269.36}$
Table 2: EE3P3D - dependence on template event count (see Sec. 3.3). Average and maximum relative errors (%) on the $12$ sequence dataset.
@@ 4.1 Dataset
Figure 3: Ground-truth frequencies of experiments in order of appearance in this work.
However, acquiring GT data this way was not feasible for all experiments. In such cases, the EE3P method [10] was used to analyse the data multiple times with varying parameter configurations to filter potential measurement outliers and obtain a GT frequency estimate as close as possible to the actual value. The frequency was then verified by manually examining the event data stream.
Event data for three experiments in this work originate from a public dataset [^5] Available at https://docs.prophesee.ai/stable/datasets.html. by Prophesee. For these experiments, we used the Event-to-Video method [^6] Available at https://bit.ly/Prophesee-E2VID. developed by Prophesee to generate a reference photo of the observed target (see our supplementary), and the EE3P method [10] with manual verification to estimate the GT frequency. All other sequences were captured as a part of this work and are publicly available. The final dataset contains sequences of period phenomena with frequencies ranging from $3.2$ Hz to 2 kHz (equivalent to 192192192192120 000120000120\,000120 000 RPM), see Fig. 3. Due to the non-constant frequencies of some periodic phenomena, we restricted our analysis to one-second segments of the event stream. This selection was based on the assumption that the change in frequency within such a time frame would be negligible and not significantly impact the accuracy of our measurements.
@@ 4.2 Measuring rotation speed
This section presents six experiments on measuring the rotation speed of objects.
4.2.1 Felt disc with a high-contrast line
For this experiment, we used a power drill to spin a white felt disc marked with a black line at $1200$ RPM (see our supplementary). Event data and tachometer readings were captured simultaneously.
All methods were expected to perform well because of the minimal noise and high contrast features present. Results confirmed this, with all methods achieving relative error lower than $0.001\%$ (see Tab. 4).
4.2.2 Fronto-parallel velcro disc
For this experiment, we used a more challenging target: a disc covered in uniform velcro material, where pattern recognition is difficult even for the human eye. A drill spun the disc at $1266$ RPM (see our supplementary).
Four out of six methods estimated the frequency despite the lack of clear features with relative error $<0.475\%$. The EB-ASM method achieved the worst accuracy (absolute error of $6228.9$ Hz, see Tab. 4). The Prophesee’s and our proposed method achieved the highest accuracy (relative error $<0.001$%), demonstrating robustness in this low-feature scenario.
@@ 4.2.2 Fronto-parallel velcro disc
Figure 4: Velcro disc with a non-frontal camera behind a glass sheet (see Sec. 4.2.3): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) a close-up photo of the target; (d) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon. Positive events are represented by white colour, negative events are bright blue.
@@
4.2.3 Velcro disc with a non-frontal camera behind a glass sheet
This experiment presented an even more challenging scenario. The velcro disc was spun at $1578$ RPM and captured at a $45\degree$ camera angle by the event camera through a glass sheet (see Fig. 4), simulating a possible industry application when the observed object is a rotating machine part and a transparent visor protects the camera and machine operators.
@@ 4.2.3 Velcro disc with a non-frontal camera behind a glass sheet
Interestingly, the Simple baseline method estimated the frequency very well ($0.38\%$ relative error). The output frequency of the FFT baseline method was roughly twice the actual frequency, likely due to the presence of centrosymmetric patterns on the velcro disc. The EB-ASM method failed to produce meaningful results. Even with the non-aligned camera axis resulting in elliptical object trajectories on the image plane, the method by Prophesee, Frequency-cam and our proposed methods produced accurate results ($<0.001$% relative error, see Tab. 4). The experiment confirms that capturing data through transparent materials does not compromise accuracy.
4.2.4 High-contrast dot
@@
For this experiment, a sequence from Prophesee’s dataset of an orbiting dark dot was used, captured by an older event camera model with a lower resolution and an older sensor generation, resulting in the sequence being stored using an older file type (RAW EVT2777See https://docs.prophesee.ai/stable/data/encoding˙formats/index.html.). The Frequency-cam could not process this file encoding, hence the lack of measurements.
The Simple baseline method did not produce any meaningful results, and the FFT baseline method estimated a frequency approximately three times higher than the GT frequency (see Tab. 3). The Prophesee’s method achieved an absolute error of $0.1$ Hz, while EB-ASM and our proposed method achieved accurate frequency estimates ($<0.001$% relative error, see Tab. 4).
4.2.5 Hand fidget spinner
This experiment used event data from Prophesee’s dataset, capturing a spinning three-blade hand fidget spinner placed on a flat surface. Because of an older file encoding, the Frequency-cam method did not provide any measurements as its dependencies do not support reading the file.
@@ 4.2.5 Hand fidget spinner
Most of the methods were unable to provide any meaningful results. The FFT baseline method ($0.3$ Hz overestimation) and our proposed method were the only two that achieved good performance despite the object being centrosymmetric (see Tab. 4). Interestingly, the method by Prophesee produced a frequency approximately four times higher than the GT frequency despite the fidget spinner having only three blades. Our proposed method achieved the best overall performance ($<0.001$% relative error, see Tab. 4).
4.2.6 Whirling Pholcus phalangioides
The sequence used for this experiment captured the whirling behaviour for evading predators used by the Pholcus phalangioides spider. This unique defence mechanism involves the spider rapidly whirling its body while its legs remain at the silk [8], creating a challenging yet intriguing scenario for measuring the rotation speed of the body of the spider.
This experiment was characterised by significant noise, dim lighting conditions, and an unstable camera, which presented one of the most challenging sequences in the dataset. A second rotating pattern was visible, caused by the shadow of the spider (see our supplementary).
Most of the methods failed to provide accurate results. The FFT baseline method estimated a frequency four times higher than the GT frequency. The method by Prophesee and Frequency-cam methods produced estimates that deviated by $1$ Hz and $2.1$ Hz, respectively, from the GT frequency. Finally, our proposed method was the only one to achieve an accurate result with a relative error of $<0.001$% (see Tab. 4).

### v2

# EEPPR: Event-based Estimation of Periodic Phenomena Rate using Correlation in 3D

@@ Abstract
We present a novel method for measuring the period of phenomena like rotation, flicker and vibration, by an event camera, a device asynchronously reporting brightness changes at independently operating pixels with high temporal resolution. The approach assumes that for a periodic phenomenon, a highly similar set of events is generated within a spatio-temporal window at a time difference corresponding to its period. The sets of similar events are detected by a correlation in the spatio-temporal event stream space. The proposed method, EEPPR, is evaluated on a dataset of $12$ sequences of periodic phenomena, i.e. flashing light and vibration, and periodic motion, e.g., rotation, ranging from $3.2$ Hz to 2 kHz (equivalent to 192192192192120 000120000120\,000120 000 RPM). EEPPR significantly outperforms published methods on this dataset, achieving the mean relative error of $0.1\%$. The dataset and codes are available on GitHub.
@@
Event camera, Frequency estimation, Periodic phenomena, Frequency of periodic phenomena.
@@ 1 Introduction
Traditional methods for measuring periodic phenomena often involve contact-based devices such as mechanical tachometers and rotary encoders. These approaches, while established, have inherent limitations: (i) physical contact with the target object can influence its movement and introduce measurement inaccuracies, (ii) contact methods may not be feasible in scenarios with delicate objects, confined spaces, or situations where interfering with the target significantly affects its movement. Laser tachometers offer a less invasive and highly accurate [^1] See https://meters.uni-trend.com/product/ut370-series/#Specifications. alternative but require reflective material (e.g., a sticker) on the target and precise aiming of the laser, as missing the reflective material pass-through results in an inaccurate measurement. Although contactless event-based methods exist, they are either handcrafted for a specific type of periodic phenomena [1] or require user supervision [13, 14].
Figure 1: EEPPR method: (i) data captured from an event camera is aggregated into a 3D array, (ii) the array is split into same-sized areas, and in each area a template depth is automatically selected, (iii) a 3D correlation of the template with the event stream is computed, (iv) a frequency is calculated from the median of time deltas measured between correlation peaks for each window, (v) the output frequency is computed as a median of measurements from all windows.
In this paper, we propose a non-contact method for measuring the properties of any periodic phenomena using an event camera. The input of our method is the event stream, and the output is a single scalar – the rate of the periodic phenomena. The method operates remotely, without requiring any modification of the observed object (e.g. a marker). The approach assumes that similar sets of pixel activations of an event camera (events) will occur at regular time intervals in any sequence capturing a periodic phenomenon. These intervals correspond to the period.
The method was evaluated on a new dataset of twelve sequences of four types of periodic phenomena: (i) flickering light, (ii) object vibration, (iii) rotation, and (iv) repetitive translational movement. A mean relative error of $0.1\%$ was achieved, which is 140$\approx$ 140$\times$ 140 × lower than the mean relative error of the best available published method (see Tab. 3). The point estimate for the period lengths is obtained by calculating a median in a spatial window for robustness. For phenomena where the time of repetition varies, the method provides the time of each period which can be further analysed.
@@ 2 Related work
Commercially available solutions for rotation speed measurement include contact and contactless methods. Contact-based devices include traditional mechanical tachometers that connect directly to a rotating shaft. These can introduce potential inaccuracies due to added mass and friction. Electrostatic sensors offer a contactless approach, detecting changes in the electromagnetic field around a rotating object (equipped with a bearing). The rotational speed is the rate of these changes [11]. Optical encoder tachometers use a photoelectric sensor to detect light passing through a rotating disc with patterned segments. The rotation speed is the frequency of light intensity changes [2]. Finally, laser tachometers utilise a laser beam that bounces off reflective markers on the rotating target. The rotation speed is the frequency of the light reflections detected by a sensor.
@@
Window size (px) $30\times 30$ $45\times 45$ $60\times 60$ $75\times 75$
Avg relative error $37.78$ $\mathbf{19.34}$ $42.96$ $487.98$
Max relative error $\mathbf{3225.00}$ $3362.50$ $7712.50$ $22737.50$
Table 1: EEPPR, average and maximum relative errors (%) as functions of window size on the dataset 5.1.
Template ev count $100-1400$ $1500-3000$
Avg relative error $308.46$ $\mathbf{5.75}$
Max relative error $22737.50$ $\mathbf{269.36}$
Table 2: EEPPR, dependence on template event count (see Sec. 4) on the dataset 5.1.
3 Baselines
A Simple baseline
was implemented to test the hypothesis that separate per-pixel events analysis is insufficient for accurate estimation of rate of complex periodic phenomena. The event camera produces a sequence of tuples $(x,y,p,t)$, where $x,y$ are spatial coordinates, $p$ is event polarity (positive – $1$ or negative – 00), and $t$ is a timestamp in microseconds of the event creation. For each pixel at position $x,y$, we construct a list of timestamps $T_{x,y,p}$ of a given polarity $p$.
@@ A Simple baseline
    Rate prediction – the final predicted rate for each window is derived from the estimated period $T$ in microseconds using the following equation $\nu(T)={10^{6}}/{T}$, where $\nu(T)$ returns rate in Hertz (Hz).
An FFT baseline
estimates per-pixel rate in four steps:
@@
    FFT computation – for every pixel and selected polarity, we compute a one-dimensional n-point Discrete Fourier Transform with the Fast Fourier Transform (FFT)  [3] with the event creation time represented by a Dirac pulse.
@@
    Median filtering – in a local 3x3 window is applied to the 2D array of estimated per-pixel rates to mitigate outliers.
@@
    Results aggregation – the mode of detected rate is returned.
4 EEPPR – The proposed method
Our method (see Fig. 1) detects periodic similarities along the time axis of event data in four steps:
@@ 4 EEPPR – The proposed method
    Correlation in 3D – a correlation in 3D of the template $\mathbf{T}_{i}$ is computed with the area $\mathbf{A}_{i}$ (see Fig. 2). The time difference between consecutive peaks in correlation responses is the estimated period of the phenomenon.
@@ 5 Experiments
In the experiments, we used the Prophesee EVK4 HD event camera with $1280\times 720$ resolution, capturing up to $1066$ million events per second [6] under light conditions ranging from $0.08$ up to 100 000100000100\,000100 000 lux [^2] Spec sheet available at https://bit.ly/Sony-Prophesee-IMX636-IMX637-specsheet. with microsecond time resolution (equivalent to 10 0001000010\,00010 000 frames per second).
@@
Since the periodic phenomena in our dataset are stationary, and there is no ego motion of the camera, we perform 3D correlation only along the time axis – “correlation in 3D”. However, in cases where the actual 3D correlation would be beneficial, it can be applied without any issues.
We evaluated the accuracy of the Simple baseline method (see Sec. 3) when using window sizes $W$ of $30\times 30$ px, $45\times 45$ px, $60\times 60$ px and $75\times 75$ px and achieved the lowest mean relative error (MRE) by setting $W$ to $45\times 45$ px. Further increasing or decreasing window size or using mode instead of the median in the temporal or spatial analysis steps of this method negatively impacted the accuracy. The FFT baseline method (see Sec. 3) achieved lower mean relative error when using timestamps of events with negative polarity (MRE of $60\%$ as opposed to $67.467\%$).
We evaluated all different configurations of parameters (window sizes $W$ of $30$, $45$, $60$ and $75$ and counts of template events $N$ of $100$, $200$, …, $3000$) for the proposed method (see Sec. 4). In all our experiments, the length of the quantisation window $t_{\text{quant}}$ is set to $100$ microseconds. When using window size $W$ set to $45\times 45$ px the method achieves the lowest MRE by a substantial margin and a low maximal relative error compared to results when using other window sizes (see Tab. 2). We noticed a significant improvement in accuracy when using counts of template events $N$ higher than $1500$ (see Tab. 2). Using 3-fold cross-validation on the introduced dataset, the best accuracy is achieved using the window size $W$ set to $45\times 45$ px and the template event count $N$ set to $1800$ events. These parameters are proposed as default parameters of the method.
We compare the results of the proposed method both with the baseline methods (see Sec. 3), and published methods – a vibration estimation method [14] developed by Prophesee, manufacturer of the used event camera, the open-source Frequency-cam method [13] and EB-ASM method [1] which we re-implemented. As the output of Frequency-cam is in the form of colourmaps, we modified it to output a reading for each pixel instead and computed their median. Since every method produces a different amount of measurements per one second of input data, we calculate the median of measurements to obtain a single value for a comparison with other methods. We followed parameter selection guidelines provided by the authors of all examined methods [14, 13, 1].
@@ 5.1 Dataset
Figure 3: Ground-truth rates of experiments in order of appearance in this work.
However, acquiring GT data this way was not feasible for all experiments. In such cases, the EE3P method [10] was used to analyse the data multiple times with varying parameter configurations to filter potential measurement outliers and obtain a GT rate estimate as close as possible to the actual value. The rate was then verified by manually examining the event data stream.
Event data for three experiments in this work originate from a public dataset [^5] Available at https://docs.prophesee.ai/stable/datasets.html. by Prophesee. For these experiments, we used the EE3P method [10] with manual verification to estimate the GT rate. All other sequences were captured as a part of this work and are publicly available. The final dataset contains sequences of period phenomena with rates ranging from $3.2$ Hz to 2 kHz (equivalent to 192192192192120 000120000120\,000120 000 RPM), see Fig. 3. Due to the non-constant rates of some periodic phenomena, we restricted our analysis to one-second segments of the event stream. This selection was based on the assumption that the change in rate within such a time frame would be negligible and not significantly impact the accuracy of our measurements.
@@ 5.2 Measuring rotation speed
Felt disc with a high-contrast line
For this experiment, we used a power drill to spin a white felt disc marked with a black line at $1200$ RPM. Event data and tachometer readings were captured simultaneously. All methods were expected to perform well because of the minimal noise and high contrast features present. Results confirmed this, with all methods achieving relative error lower than $0.001\%$ (see Tab. 3).
Fronto-parallel velcro disc
For this experiment, we used a more challenging target: a disc covered in uniform velcro material, where pattern recognition is difficult even for the human eye. A drill spun the disc at $1266$ RPM. Four out of six methods estimated the rate despite the lack of clear features with relative error $<0.475\%$. The EB-ASM method achieved the worst accuracy (absolute error of $6228.9$ Hz, see Tab. 3). The Prophesee’s and our proposed method achieved the highest accuracy (relative error $<0.001$%), demonstrating robustness in this low-feature scenario.
@@
Figure 4: Velcro disc with a non-frontal camera behind a glass sheet (see Sec. 5.2): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) a close-up photo of the target; (d) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon. Positive events are represented by white colour, negative events are bright blue.
@@
Velcro disc with a non-frontal camera behind a glass sheet
@@
This experiment presented an even more challenging scenario. The velcro disc was spun at $1578$ RPM and captured at a $45\degree$ camera angle by the event camera through a glass sheet (see Fig. 4), simulating a possible industry application when the observed object is a rotating machine part and a transparent visor protects the camera and machine operators. Interestingly, the Simple baseline method estimated the rate very well ($0.38\%$ relative error). The output frequency of the FFT baseline method was roughly twice the actual rate, likely due to the presence of centrosymmetric patterns on the velcro disc. The EB-ASM method failed to produce meaningful results. Even with the non-aligned camera axis resulting in elliptical object trajectories on the image plane, the method by Prophesee, Frequency-cam and our proposed methods produced accurate results ($<0.001$% relative error, see Tab. 3). The experiment confirms that capturing data through transparent materials does not compromise accuracy.
High-contrast dot
@@
A sequence from Prophesee’s dataset of an orbiting dark dot was used, captured by an older event camera model with a lower resolution and an older sensor generation, resulting in the sequence being stored using an older file type (RAW EVT2666See https://docs.prophesee.ai/stable/data/encoding˙formats/index.html.). The Frequency-cam could not process this file encoding, hence the lack of measurements. The Simple baseline method did not produce any meaningful results, and the FFT baseline method estimated a frequency approximately three times higher than the GT rate. The Prophesee’s method achieved an absolute error of $0.1$ Hz, while EB-ASM and our proposed method achieved accurate rate estimates ($<0.001$% relative error, see Tab. 3).
Hand fidget spinner
@@
This experiment used event data from Prophesee’s dataset, capturing a spinning three-blade hand fidget spinner placed on a flat surface. Because of an older file encoding, the Frequency-cam method did not provide any measurements as its dependencies do not support reading the file. Most of the methods were unable to provide any meaningful results. The FFT baseline method ($0.3$ Hz overestimation) and our proposed method were the only two that achieved good performance despite the object being centrosymmetric. Interestingly, the method by Prophesee produced a rate approximately four times higher than the GT rate despite the fidget spinner having only three blades. Our proposed method achieved the best overall performance ($<0.001$% relative error, see Tab. 3).
Whirling Pholcus phalangioides
The sequence used for this experiment captured the whirling behaviour for evading predators used by the Pholcus phalangioides spider. This unique defence mechanism involves the spider rapidly whirling its body while its legs remain at the silk [8], creating a challenging yet intriguing scenario for measuring the rotation speed of the body of the spider. This experiment was characterised by significant noise, dim lighting conditions, and an unstable camera, which presented one of the most challenging sequences in the dataset. A second rotating pattern was visible, caused by the shadow of the spider. Most of the methods failed to provide accurate results. The FFT baseline method estimated a frequency four times higher than the GT rate. The method by Prophesee and Frequency-cam methods produced estimates that deviated by $1$ Hz and $2.1$ Hz, respectively, from the GT rate. Finally, our proposed method was the only one to achieve an accurate result with a relative error of $<0.001$% (see Tab. 3).